Train A and B are traveking the same directon on parrallel tracks.Train A is traveling 80mph and Train B is traveling 90 mph. Train A passes a station at 4:20 and Train B passes the same stattion at 4:32. How long will it take for train B to catch up with train A? What time will it be?
Answer by nerdybill(7384)

Train A and B are traveking the same directon on parrallel tracks.Train A is traveling 80mph and Train B is traveling 90 mph. Train A passes a station at 4:20 and Train B passes the same stattion at 4:32. How long will it take for train B to catch up with train A? What time will it be?
.
Essentially, Train A (the slower train) has a 12 minute head start. 12/60 is .2 hours head start.
.
Let x = hours it takes for train B to catch up
then
(x+.2) = time Train A traveled
.
For Train B to catch up with Train A, the distance traveled since the station MUST be the same: d=rt (distance=rate*time)
80(x+.2) = 90x
80x + 16 = 90x
16 = 10x
1.6 hrs = x (answer to first question)
.
4:20 + 1.6 hrs
.6 hrs is .6*60= 36min
therefore, 4:20 + 1:36 = 5:56